# LLM-based Few-Shot Early Rumor Detection with Imitation Agent

### v1

@@ 1. Introduction
EARD aims to automatically determine an early time point in a sequence of social posts related to an unverified claim, at which the prediction regarding whether that claim is a rumor shall be accurate (zeng-gao-2022-early). Specifically, the model continuously monitors a stream of posts, and at each time step it must assess whether the current moment is suitable for making an accurate prediction (i.e., classifying the claim as a rumor or non-rumor) based on the observed social posts. If it determines that such a decision can be made, the model outputs its prediction and terminates further observation; otherwise, it continues monitoring the stream.
@@
The EARD task is challenging yet has received limited attention. Existing EARD methods typically rely on recurrent neural networks (RNNs) for the rumor detection module, and the early time point prediction module often employs deep reinforcement learning (zhou-etal-2019-early), fixed probability thresholds (song2019ced), or the neural Hawkes Process (zeng-gao-2022-early). These methods tightly couple the two modules for joint training in a recurrent manner, necessitating extensive annotated data for effective model training while lacking the flexibility to easily replace either module with a more powerful alternative.
@@
New events on social media constantly emerge, while there could be rarely or even no annotated claims for them and each claim tends to have limited number of relevant posts available in the early stage of propagation. Meanwhile, the evolving nature of social conversations about rumorous events requires advanced text understanding capabilities for a detector. Annotating training data under such dynamics is very costly (saakyan-etal-2021-covid). Moreover, waiting for lots of posts to accumulate before making decision (e.g., for a intervention) can inadvertently allow rumors to spread even further. This dilemma highlights the critical need for effective EARD methods in data-limited scenarios, a challenge that existing approaches struggle to address.
@@
Recently, Large Language Models (LLMs) have demonstrated promising capabilities in understanding social media text with limited data across various tasks (10.1145/3543873.3587605; zhang-etal-2024-sentiment; yang-etal-2024-reinforcement; Lan_Gao_Jin_Li_2024; 10459901; roy-etal-2023-probing; liu2024largelanguagemodelsdetect). However, LLMs face significant limitations when applied to the EARD task, as they are not inherently well-suited for handling time series data or representing sequential dependencies of such data (LLM-TS) while training LLMs for dealing with such data requires intensive resources. Additionally, in the context of EARD, the cost of inference can become prohibitively high because the model must generate predictions at every time step along the sequence of posts, with the context length increasing over time, further compounding computational overhead and latency. This raises an important question: Given the strong language understanding capabilities of LLMs, how can we efficiently and effectively utilize them for few-shot EARD?
@@
We first model EARD as a Markov Decision Process (MDP), where the observed posts represent the state, the action involves deciding whether to trigger the LLM for rumor detection. The agent aims to maximize expected returns, reflecting its ability to make early and accurate detection. However, the expected return depends on an effective reward function that can be highly challenging to design in a complex, real-world setting (ng2000algorithms). In EARD, reward for determining the optimal early time point is especially difficult to quantify, relying on the sufficiency of observed information and the capability of the detector. To address this, we adopt imitation learning (IL), which enables the agent to learn optimal policy from a few expert trajectories, ultimately imitating the expert’s decision-making process. IL has been successfully applied to many sequential decision-making problems (kuefler2017imitating; hussein2017imitation; ingimundardottir2018discovering; fang2019survey), and in our case, it allows us to bypass the need for explicitly defining complex reward functions. Specifically, we design three types of strategies to curate expert trajectories based on the LLM’s predictions at each time step along the sequence: the first two focus on state-action pairs that lead to early, stable, and correct predictions, and the third captures the state-action pairs where the LLM fails to make a correct prediction even when considering all available posts. Consequently, we obtain three sets of state-action pairs, and the distribution of each set is referred to as the occupancy measure, which accounts for both expert’s policy and sequence dynamics. Our goal is to minimize the distance between the occupancy measure of the agent and those of the first two experts, while maximizing the divergence from the third expert. This is to ensure that the agent learns an optimal policy for better generalization by effectively imitating correct prediction patterns and avoiding the behaviors represented by incorrect predictions.
@@
  • Extensive experiments on four real-world rumor detection datasets demonstrate the effectiveness of our proposed method across various LLMs, including Mistral (Mistral), Llama 3 (llama3), ChatGPT (ChatGPT). Our approach also outperforms existing EARD methods in terms of detection accuracy and earliness metrics.
@@ 2. Related Work
Many studies in rumor detection assert that their models can be generally adopted for early detection by simply feeding them data observed up to a predetermined checkpoint that could be either a specific time point or number of observed posts (ma2016detecting; yu2017convolutional; ma2017detect; ma-etal-2018-rumor; guo2018rumor; bian2020rumor; Lin_Yi_Ma_Jiang_Luo_Shi_Liu_2023). Yet, such methods do not consider how to determine an optimal detection point. Typically, the checkpoints are chosen manually and coarse-grained. It is infeasible to determine an optimal early detection point among these checkpoints, since the decision requires referring to data points or model outputs beyond the current decision point (e.g., for the stability of prediction), thus delaying detection. Some other methods claim to be designed for early detection by integrating temporal factor into their classifier (zhao2015enquiring; nguyen2017early; wu2017gleaning; xia-etal-2020-state), but they still lack mechanisms to enforce earliness and cannot automatically establish an optimal detection point.
There are several full-shot EARD methods that is capable of automatically deciding whether to stop or continue at a checkpoint. A method called ERD (zhou-etal-2019-early) used deep reinforcement learning with a manually designed reward function to encourage the model to focus on early time intervals during training.  song2019ced proposed another EARD method named Credible Detection Point (CED), which used a fixed probability threshold to determine whether to halt the detection process based on the credibility of the current prediction. Later, HEARD (zeng-gao-2022-early) highlighted the instability and low confidence of these models, noticing that they fail to account for the uncertainty of future predictions during training. HEARD addressed this by employing the neural Hawkes Process (mei-etal-2017-The) to construct a detection stability distribution based on a sequence of prior and current predictions, which allows the model to automatically determine a detection point at which predictions in future time steps remain unchanged. However, such cautiousness can lead to delayed detection as the decision is pending for achieving stabilization, while the noisy posts during rumor propagation can easily disturb the stabilization process. Also, these methods are full-shot, requiring extensive annotated data for effective training, which makes them unsuitable for data-limited scenarios. In contrast, our few-shot method is specifically designed for the situation where only a few labeled instances are available, thereby aligning with real-world data limitations especially in EARD.
@@ 4.2. Imitation Learning
In the MDP formulation for EARD, the agent aims to find an optimal policy $\pi^{*}$ that maximizes the expected return, which is fundamentally influenced by the reward function measuring how early and accurate detection can be achieved. Reinforcement Learning (RL) methods typically rely on well-defined, observable reward function to guide the agent’s actions. However, creating an effective reward function can be extremely challenging in many real-world scenarios (ng2000algorithms), including EARD.
In EARD, determining the early time point relies on both the observed information and the detector’s capability, both of which are difficult to quantify. Manually specified reward functions may be sparse or misaligned with true task objectives, making it difficult for the agent to learn an optimal policy and potentially leading to unintended behaviors that might cause delayed predictions (zeng-gao-2022-early). Learning a reward function can be costly, requiring large labeled datasets and substantial computational resources to train predefined neural networks (ho2016generative; RLHF-cons).
To address these challenges, we utilize imitation learning (IL) to train an autonomous agent, offering a practical alternative that has been successfully applied in many sequential decision-making problems (kuefler2017imitating; hussein2017imitation; ingimundardottir2018discovering; fang2019survey). IL is particularly advantageous in environments where designing an explicit reward function is difficult or infeasible (ng2000algorithms), as it enables the agent to learn from expert demonstrations without the need for an explicitly predefined reward function. IL can achieve good performance with fewer data samples (bain1995framework), making it more efficient compared to RL that often requires extensive exploration. Moreover, by directly mimicking the decision-making process of the expert, IL typically converges faster to effective policies than RL which relies on trial-and-error exploration to discover optimal actions (ross2011reduction). In the IL setting, we are provided with a limited number of expert trajectories from experts. These trajectories consist of state-action pairs that illustrate the actions taken by the expert in response to various states in the environment. Our goal is to learn a policy for the agent that imitates the expert’s behavior.
@@ 4.2.1. Trajectories Generation
Conservative Expert (CE): Inspired by zeng-gao-2022-early, this expert determines the early time point $t_{i^{*}}$, at which the prediction for a claim is accurate and remains unchanged over time. This means there are no reversals in the prediction after $t_{i^{*}}$; formally, $t_{i^{*}}=\text{min}\{t_{k}\;|\;\hat{y}_{k}=\hat{y}_{j}=y,\text{for all}\;j,\;k<j\leq|M|,k\in\{0,\ldots,|M|\}\}$. The stop action is taken when predictions are stable without further changes. However, this cautious approach can delay detection, as it is always likely that the model might be misled to change its prediction by irrelevant or distracting posts during rumor propagation.
@@ 4.2.2. Objective Function
With our trajectory generation strategies, we obtain three sets of expert trajectories $\Phi_{E}=\{\tau_{E}\sim\pi_{E}\}$ for training, where $E=\{c,e,m\}$ correspond to CE, EAE, and ME, respectively. These trajectories consist of state-action pairs that demonstrate the expert policy $\pi_{E}$ in the given environment. The agent is expected to learn an optimal policy that synthesizes the policies of different experts, i.e., learning from desirable expert behaviors while suppressing those from misleading ones. To achieve this, we employ generative adversarial imitation learning (ho2016generative), a widely used imitation learning algorithm that frames the imitation learning problem as matching the occupancy measures between the expert and agent policies. The occupancy measure reflects the distribution of state-action pairs and accounts for both the policy and the environment dynamics. Specifically, the occupancy measure of $\pi$ is denoted as $\rho_{\pi}\in\Gamma$, defined as $\rho_{\pi}(s,a)=\pi(a|s)\sum_{t=0}^{\infty}\bar{\gamma}^{t}\mathcal{P}(s_{t}=s|\pi)$ (puterman2014markov), where $\Gamma$ is the set of occupancy measures associated with all stationary stochastic policies $\Pi$.
Since there is an one-to-one correspondence between $\Pi$ and $\Gamma$ (syed2008apprenticeship), learning a policy $\pi$ can be framed as a matching problem between the occupancy measure of the agent policy $\rho_{\pi}$ and that of the expert policy $\rho_{\pi_{E}}$ (ho2016generative). Therefore, in our case, the agent learns a policy $\pi$ whose occupancy measure $\rho_{\pi}$ minimizes the divergence between $\rho_{\pi_{c}}$ and $\rho_{\pi_{e}}$, while maximizing the divergence from $\rho_{\pi_{m}}$. The objective function is defined as follows:
@@
Following ho2016generative, $\psi^{*}$ is defined as:
@@
where $D\in(0,1)^{\mathcal{S}\times\mathcal{A}}$ is the discriminative classifiers responsible for distinguishing between state-action pairs of agent policy and those of the expert. This equation represents the optimal negative log loss of this binary classification problem of differentiating agent and expert trajectories. This optimal loss is the Jensen-Shannon divergence $JSD(\rho_{\pi},\rho_{\pi_{E}})$ (ho2016generative), which provides a distance metric between occupancy measures. The agent aims to minimize the divergence between its state-action distribution and those of CE and EAE, while maximizing the divergence from ME.
@@ 4.2.3. Training
To solve the objective function in Eq. (1), following ho2016generative, our agent consists of three neural networks: a policy network $\pi_{\theta}$ parameterized by $\theta$, a discriminator network $D_{\omega}$ parameterized by $\omega$, and a value network $V_{\phi}$ parameterized by $\phi$. During training, the agent alternates between collecting trajectories using the current policy and updating the policy via Proximal Policy Optimization (PPO) method (schulman2017proximal). The discriminator distinguishes expert trajectories from agent-generated ones and provides an implicit reward signal. Since PPO method requires estimating the advantage of each action (schulman2017proximal), the value network serves as the critic and estimates the state-value function $V_{\phi}(s)$, stabilizing policy updates.
@@
Our approach aims to compute a saddle point $(\theta,\omega)$ for the above objective function, by updating the parameters of policy and discriminator sequentially (achiam2017constrained; shao2024imitating; satija2020constrained):
@@
We use Adam optimizer (kingma2014adam) to update the variable $\omega$, targeting the maximization of Eq. (3) with respect to the discriminator $D_{\omega}$.
@@
where $\epsilon$ is a small hyperparameter that limits how far the new policy $\pi_{\theta}$ can deviate from the old policy $\pi_{\theta_{k}}$, and $A^{\pi_{\theta_{k}}}$ is the Advantage that measures the difference between the Q-value and the value function $V_{\phi}$ parameterized by $\phi$. Since no explicit reward function is available to compute the Advantage $A^{\pi_{\theta_{k}}}$, we use the output of the discriminator, $\log D_{\omega}(s,a)$, as the implicit reward signal. We then apply Generalized Advantage Estimation (GAE) (schulman2015high) to compute $A^{\pi_{\theta_{k}}}(s,a)$.
@@
where $\hat{R}_{t}$ represents the reward-to-go, which is also computed using the GAE method (schulman2015high).

### v2

@@ 1. Introduction
EARD aims to automatically determine an early time point in a sequence of social posts related to an unverified claim, at which the prediction regarding whether that claim is a rumor shall be accurate (Zeng and Gao, 2022). Specifically, the model continuously monitors a stream of posts, and at each time step it must assess whether the current moment is suitable for making an accurate prediction (i.e., classifying the claim as a rumor or non-rumor) based on the observed social posts. If it determines that such a decision can be made, the model outputs its prediction and terminates further observation; otherwise, it continues monitoring the stream.
@@
The EARD task is challenging yet has received limited attention. Existing EARD methods typically rely on recurrent neural networks (RNNs) for the rumor detection module, and the early time point prediction module often employs deep reinforcement learning (Zhou et al., 2019a), fixed probability thresholds (Song et al., 2019), or the neural Hawkes Process (Zeng and Gao, 2022). These methods tightly couple the two modules for joint training in a recurrent manner, necessitating extensive annotated data for effective model training while lacking the flexibility to easily replace either module with a more powerful alternative.
@@
New events on social media constantly emerge, while there could be rarely or even no annotated claims for them and each claim tends to have limited number of relevant posts available in the early stage of propagation. Meanwhile, the evolving nature of social conversations about rumorous events requires advanced text understanding capabilities for a detector. Annotating training data under such dynamics is very costly (Saakyan et al., 2021). Moreover, waiting for lots of posts to accumulate before making decision (e.g., for a intervention) can inadvertently allow rumors to spread even further. This dilemma highlights the critical need for effective EARD methods in data-limited scenarios, a challenge that existing approaches struggle to address.
@@
Recently, Large Language Models (LLMs) have demonstrated promising capabilities in understanding social media text with limited data across various tasks (Deng et al., 2023; Zhang et al., 2024; Yang et al., 2024; Lan et al., 2024; Guo et al., 2023; Roy et al., 2023; Liu et al., 2024). However, LLMs face significant limitations when applied to the EARD task, as they are not inherently well-suited for handling time series data or representing sequential dependencies of such data (Tan et al., 2024) while training LLMs for dealing with such data requires intensive resources. Additionally, in the context of EARD, the cost of inference can become prohibitively high because the model must generate predictions at every time step along the sequence of posts, with the context length increasing over time, further compounding computational overhead and latency. This raises an important question: Given the strong language understanding capabilities of LLMs, how can we efficiently and effectively utilize them for few-shot EARD?
@@
We first model EARD as a Markov Decision Process (MDP), where the observed posts represent the state, the action involves deciding whether to trigger the LLM for rumor detection. The agent aims to maximize expected returns, reflecting its ability to make early and accurate detection. However, the expected return depends on an effective reward function that can be highly challenging to design in a complex, real-world setting (Ng et al., 2000). In EARD, reward for determining the optimal early time point is especially difficult to quantify, relying on the sufficiency of observed information and the capability of the detector. To address this, we adopt imitation learning (IL), which enables the agent to learn optimal policy from a few expert trajectories, ultimately imitating the expert’s decision-making process. IL has been successfully applied to many sequential decision-making problems (Kuefler et al., 2017; Hussein et al., 2017; Ingimundardottir and Runarsson, 2018; Fang et al., 2019), and in our case, it allows us to bypass the need for explicitly defining complex reward functions. Specifically, we design three types of strategies to curate expert trajectories based on the LLM’s predictions at each time step along the sequence: the first two focus on state-action pairs that lead to early, stable, and correct predictions, and the third captures the state-action pairs where the LLM fails to make a correct prediction even when considering all available posts. Consequently, we obtain three sets of state-action pairs, and the distribution of each set is referred to as the occupancy measure, which accounts for both expert’s policy and sequence dynamics. Our goal is to minimize the distance between the occupancy measure of the agent and those of the first two experts, while maximizing the divergence from the third expert. This is to ensure that the agent learns an optimal policy for better generalization by effectively imitating correct prediction patterns and avoiding the behaviors represented by incorrect predictions.
@@
  • Extensive experiments on four real-world rumor detection datasets demonstrate the effectiveness of our proposed method across various LLMs, including Mistral (Jiang et al., 2023), Llama 3 (Dubey et al., 2024), ChatGPT (OpenAI, 2024). Our approach also outperforms existing EARD methods in terms of detection accuracy and earliness metrics.
@@ 2. Related Work
Many studies in rumor detection assert that their models can be generally adopted for early detection by simply feeding them data observed up to a predetermined checkpoint that could be either a specific time point or number of observed posts (Ma et al., 2016; Yu et al., 2017; Ma et al., 2017, 2018; Guo et al., 2018; Bian et al., 2020; Lin et al., 2023). Yet, such methods do not consider how to determine an optimal detection point. Typically, the checkpoints are chosen manually and coarse-grained. It is infeasible to determine an optimal early detection point among these checkpoints, since the decision requires referring to data points or model outputs beyond the current decision point (e.g., for the stability of prediction), thus delaying detection. Some other methods claim to be designed for early detection by integrating temporal factor into their classifier (Zhao et al., 2015; Nguyen et al., 2017; Wu et al., 2017; Xia et al., 2020), but they still lack mechanisms to enforce earliness and cannot automatically establish an optimal detection point.
There are several full-shot EARD methods that is capable of automatically deciding whether to stop or continue at a checkpoint. A method called ERD (Zhou et al., 2019a) used deep reinforcement learning with a manually designed reward function to encourage the model to focus on early time intervals during training.  Song et al. (2019) proposed another EARD method named Credible Detection Point (CED), which used a fixed probability threshold to determine whether to halt the detection process based on the credibility of the current prediction. Later, HEARD (Zeng and Gao, 2022) highlighted the instability and low confidence of these models, noticing that they fail to account for the uncertainty of future predictions during training. HEARD addressed this by employing the neural Hawkes Process (Mei and Eisner, 2017) to construct a detection stability distribution based on a sequence of prior and current predictions, which allows the model to automatically determine a detection point at which predictions in future time steps remain unchanged. However, such cautiousness can lead to delayed detection as the decision is pending for achieving stabilization, while the noisy posts during rumor propagation can easily disturb the stabilization process. Also, these methods are full-shot, requiring extensive annotated data for effective training, which makes them unsuitable for data-limited scenarios. In contrast, our few-shot method is specifically designed for the situation where only a few labeled instances are available, thereby aligning with real-world data limitations especially in EARD.
@@ 4.2. Imitation Learning
In the MDP formulation for EARD, the agent aims to find an optimal policy $\pi^{*}$ that maximizes the expected return, which is fundamentally influenced by the reward function measuring how early and accurate detection can be achieved. Reinforcement Learning (RL) methods typically rely on well-defined, observable reward function to guide the agent’s actions. However, creating an effective reward function can be extremely challenging in many real-world scenarios (Ng et al., 2000), including EARD.
In EARD, determining the early time point relies on both the observed information and the detector’s capability, both of which are difficult to quantify. Manually specified reward functions may be sparse or misaligned with true task objectives, making it difficult for the agent to learn an optimal policy and potentially leading to unintended behaviors that might cause delayed predictions (Zeng and Gao, 2022). Learning a reward function can be costly, requiring large labeled datasets and substantial computational resources to train predefined neural networks (Ho and Ermon, 2016; Casper et al., 2023).
To address these challenges, we utilize imitation learning (IL) to train an autonomous agent, offering a practical alternative that has been successfully applied in many sequential decision-making problems (Kuefler et al., 2017; Hussein et al., 2017; Ingimundardottir and Runarsson, 2018; Fang et al., 2019). IL is particularly advantageous in environments where designing an explicit reward function is difficult or infeasible (Ng et al., 2000), as it enables the agent to learn from expert demonstrations without the need for an explicitly predefined reward function. IL can achieve good performance with fewer data samples (Bain and Sammut, 1995), making it more efficient compared to RL that often requires extensive exploration. Moreover, by directly mimicking the decision-making process of the expert, IL typically converges faster to effective policies than RL which relies on trial-and-error exploration to discover optimal actions (Ross et al., 2011). In the IL setting, we are provided with a limited number of expert trajectories from experts. These trajectories consist of state-action pairs that illustrate the actions taken by the expert in response to various states in the environment. Our goal is to learn a policy for the agent that imitates the expert’s behavior.
@@ 4.2.1. Trajectories Generation
Conservative Expert (CE): Inspired by Zeng and Gao (2022), this expert determines the early time point $t_{i^{*}}$, at which the prediction for a claim is accurate and remains unchanged over time. This means there are no reversals in the prediction after $t_{i^{*}}$; formally, $t_{i^{*}}=\text{min}\{t_{k}\;|\;\hat{y}_{k}=\hat{y}_{j}=y,\text{for all}\;j,\;k<j\leq|M|,k\in\{0,\ldots,|M|\}\}$. The stop action is taken when predictions are stable without further changes. However, this cautious approach can delay detection, as it is always likely that the model might be misled to change its prediction by irrelevant or distracting posts during rumor propagation.
@@ 4.2.2. Objective Function
With our trajectory generation strategies, we obtain three sets of expert trajectories $\Phi_{E}=\{\tau_{E}\sim\pi_{E}\}$ for training, where $E=\{c,e,m\}$ correspond to CE, EAE, and ME, respectively. These trajectories consist of state-action pairs that demonstrate the expert policy $\pi_{E}$ in the given environment. The agent is expected to learn an optimal policy that synthesizes the policies of different experts, i.e., learning from desirable expert behaviors while suppressing those from misleading ones. To achieve this, we employ generative adversarial imitation learning (Ho and Ermon, 2016), a widely used imitation learning algorithm that frames the imitation learning problem as matching the occupancy measures between the expert and agent policies. The occupancy measure reflects the distribution of state-action pairs and accounts for both the policy and the environment dynamics. Specifically, the occupancy measure of $\pi$ is denoted as $\rho_{\pi}\in\Gamma$, defined as $\rho_{\pi}(s,a)=\pi(a|s)\sum_{t=0}^{\infty}\bar{\gamma}^{t}\mathcal{P}(s_{t}=s|\pi)$ (Puterman, 2014), where $\Gamma$ is the set of occupancy measures associated with all stationary stochastic policies $\Pi$.
Since there is an one-to-one correspondence between $\Pi$ and $\Gamma$ (Syed et al., 2008), learning a policy $\pi$ can be framed as a matching problem between the occupancy measure of the agent policy $\rho_{\pi}$ and that of the expert policy $\rho_{\pi_{E}}$ (Ho and Ermon, 2016). Therefore, in our case, the agent learns a policy $\pi$ whose occupancy measure $\rho_{\pi}$ minimizes the divergence between $\rho_{\pi_{c}}$ and $\rho_{\pi_{e}}$, while maximizing the divergence from $\rho_{\pi_{m}}$. The objective function is defined as follows:
@@
Following Ho and Ermon (2016), $\psi^{*}$ is defined as:
@@
where $D\in(0,1)^{\mathcal{S}\times\mathcal{A}}$ is the discriminative classifiers responsible for distinguishing between state-action pairs of agent policy and those of the expert. This equation represents the optimal negative log loss of this binary classification problem of differentiating agent and expert trajectories. This optimal loss is the Jensen-Shannon divergence $JSD(\rho_{\pi},\rho_{\pi_{E}})$ (Ho and Ermon, 2016), which provides a distance metric between occupancy measures. The agent aims to minimize the divergence between its state-action distribution and those of CE and EAE, while maximizing the divergence from ME.
@@ 4.2.3. Training
To solve the objective function in Eq. (1), following Ho and Ermon (2016), our agent consists of three neural networks: a policy network $\pi_{\theta}$ parameterized by $\theta$, a discriminator network $D_{\omega}$ parameterized by $\omega$, and a value network $V_{\phi}$ parameterized by $\phi$. During training, the agent alternates between collecting trajectories using the current policy and updating the policy via Proximal Policy Optimization (PPO) method (Schulman et al., 2017). The discriminator distinguishes expert trajectories from agent-generated ones and provides an implicit reward signal. Since PPO method requires estimating the advantage of each action (Schulman et al., 2017), the value network serves as the critic and estimates the state-value function $V_{\phi}(s)$, stabilizing policy updates.
@@
Our approach aims to compute a saddle point $(\theta,\omega)$ for the above objective function, by updating the parameters of policy and discriminator sequentially (Achiam et al., 2017; Shao et al., 2024; Satija et al., 2020):
@@
We use Adam optimizer (Kingma and Ba, 2015) to update the variable $\omega$, targeting the maximization of Eq. (3) with respect to the discriminator $D_{\omega}$.
@@
where $\epsilon$ is a small hyperparameter that limits how far the new policy $\pi_{\theta}$ can deviate from the old policy $\pi_{\theta_{k}}$, and $A^{\pi_{\theta_{k}}}$ is the Advantage that measures the difference between the Q-value and the value function $V_{\phi}$ parameterized by $\phi$. Since no explicit reward function is available to compute the Advantage $A^{\pi_{\theta_{k}}}$, we use the output of the discriminator, $\log D_{\omega}(s,a)$, as the implicit reward signal. We then apply Generalized Advantage Estimation (GAE) (Schulman et al., 2015) to compute $A^{\pi_{\theta_{k}}}(s,a)$.
@@
where $\hat{R}_{t}$ represents the reward-to-go, which is also computed using the GAE method (Schulman et al., 2015).